.1.1 Use Euler’s method with step size h = 0.1 to approximate the solution to the initial valueproblem: y'=x y, y(1)=4, atthepointsx=1.1,1.2and1.3,correcttofivedecimal

places.

2.1.2 If the analytical solution to the initial value problem in (2.1.1) is

(4) y = 1 (x2 + 7), determine

4

the %Error in the numerical method where x =1.3

2.2 A chicken cools down from 100 °C to 60 °C within 10 


1
Expert's answer
2021-12-14T15:58:05-0500

2.1.1

"y'=x\sqrt y,y(1)=4"


Euler method

"y_1=y_0+hf(x_0,y_0)=4+0.1f(1,4)=4.2"

 "y_2=y_1+hf(x_1,y_1)=4.2+0.1f(1.1,4.2)=4.42543"

"y_3=y_2+hf(x_2,y_2)=4.42543+0.1f(1.2,4.42543)=4.67787"

"y(1.3)=4.67787"


2.1.2

"\sqrt y=(x^2+7)/4"

"y(1.3)=(1.3^2+7)^2\/16=4.71976"


error = "\frac{4.71976-4.67787}{4.71976}=0.0089=0.89\%"


2.2

A chicken cools down from 100 °C to 60 °C within 10 minutes of being placed in a stream of air

at "20\degree C"


by Newton's Law of Cooling:

rate of cooling:

"\frac{dT}{dt}=-k(T-T_{env})"

where k is constant,

Tenv is temperature of the environment


then:

"\frac{dT}{dt}=-k(T-20)"


"T=ce^{-kt}-20"

"T(0)=c-20=100"

"c=120"

"T(10)=120e^{-10k}-20=60"

"k=-ln(80\/120)\/10=0.04"


"T(t)=120e^{-0.04t}-20"